OpenStudy (anonymous):

How many solutions exist for the system of equations? (Use the graphing calculator tool to help solve the problem.) –2y + 14x = 4 y = 7x + 14

OpenStudy (anonymous):

a. one solution: approximately (–0.9, 8) b. one solution: approximately (8, –0.9) c. no solution d. infinite number of solutions

OpenStudy (anonymous):

what english u have

OpenStudy (anonymous):

what number is that i think i did that

OpenStudy (anonymous):

it's on a quiz

OpenStudy (anonymous):

when can u help on english1 because i need help 02.06A The Heart of the Journey

OpenStudy (perl):

–2y + 14x = 4 14x - 4 = 2y 7x - 2 = y y = 7x + 14 y = 7x - 2 ------------------ 0 = 14 , no solution